Circular No. 6938

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980613 M. J. S. Smith, BeppoSAX Science Operation Center (SOC), Rome,

and Space Research Organization of the Netherlands (SRON); R. Ricci, SOC; J. in ’t Zand, SRON; M. Feroci, L. Piro, and P. Soffitta, Istituto di Astrofisica Spaziale, CNR, Rome; and N. Masetti and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, communicate: “A weak γ-ray burst was observed by the BeppoSAX Gamma Ray Burst Monitor on June 13.20215 UT. The preliminary analysis shows a duration of ∼ 50 s and a peak intensity of 250 counts/s in the energy band 40–700 keV. The burst was also detected in Wide Field Camera (WFC) unit 2 with comparable duration and a peak intensity of ∼ 0.6 ± 0.1 Crab (2–25 keV). The preliminary position as derived from the WFC image is α = 10^{h}17^{m}44^{s}, δ = +71^{o}29′.9 (equinox 2000.0), with an error radius of 4′. A follow-up observation with BeppoSAX narrow-field instruments has been in progress since June 13.56. Observations in all bands are solicited.”

XTE J1748–288 M. P. Rupen, R. M. Hjellming, National Radio Astronomy Observa-

tory (NRAO); and A. J. Mioduszewski, Joint Institute for VLBI in Europe and NRAO, report: “VLA imaging of the highly variable radio counterpart of XTE J1748–288 (IAUC 6934, 6937) show that, although preliminary analysis indicates it was an unresolved source with a size < 0′′.1 on June 10.19 UT, on June 14.31 it was extended east-west with size ∼ 0′′.25, indicating proper motions of ≥ 20–40 mas/day, depending upon whether it was two-sided or one-sided jet ejection. June 15.2 observations confirmed this result. Preliminary fluxes for June 14.31 are 590, 410, 330, 280, and 200 mJy at 1.46, 4.86, 8.46, 14.9, and 22.5 GHz, respectively.”

V4334 SAGITTARII H. U. Kaeufl, European Southern Observatory (ESO); and B. Steck-

lum, Thüringer Landessternwarte, Tautenburg, write: “Using the ESO 3.6m telescope (+ thermal infrared instrument TIMMI) on June 12, we observed V4334 Sgr (cf. IAUC 6601, 6896) in the atmospheric N band in two filters. The observed fluxes were 5.7 Jy in a silicon carbide filter (10.3–13.0 μm) and 8.2 Jy in the 9.7–11.3-μm filter. The object is of stellar appearance. The upper limit for a potential apparent spatial extent of the infrared emitting region of V4334 Sgr is 1′′.0.”

1998 June 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green